OpenStudy (anonymous):

6m – 2 = 10 – 6(2 – m)

OpenStudy (anonymous):

Do you want to solve the equation ?

OpenStudy (anonymous):

yes do I distribute 1st

OpenStudy (anonymous):

Yes, and then ? tell us what you get !

OpenStudy (anonymous):

6m-2=10-12m-6m

OpenStudy (anonymous):

NO ! There is an error with the last sign It is (+6m) not (-6m) because you have \[-6\times(-m)\]

OpenStudy (anonymous):

o

OpenStudy (anonymous):

So, when you distribute the brackets you get : \[6m – 2 = 10 -12 +6m\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is it m=4

OpenStudy (anonymous):

You observe that 6m is in the two sides of the equality, so we delete it ! It remains : \[-2=10-12\] which is true ! So, The set of solution is the set of all real numbers !

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

-2=-2

OpenStudy (anonymous):

Yes, you get -2=-2, this means that the equation is always true for all value of m So : The set of solution is the set of all real numbers ! In other words! : \[S=\mathbb R\]